how many liters of 9% salt solution and 15% salt solution are needed to make 18 liters of 10% salt solution
Answer by Alan3354(69443)

how many liters of 9% salt solution and 15% salt solution are needed to make 18 liters of 10% salt solution
---------
n = 9%
f = 15%
------
n+f = 18 liters
0.09n + 0.15f = 0.1*18 (total of salt)
9n + 15f = 180
n = 18-f
Sub for n
9(18-f) + 15f = 180
162 - 9f + 15f = 180
6f = 18
f = 3 liters
n = 15 liters